THE PERMANENT ELECTRIC DIPOLE MOMENTS OF IRON MONOCARBIDE

\underline{TIMOTHY C. STEIMLE} and WILTON L. VIRGO, Department of Chemistry and Biochemistry, Arizona State University, Tempe, AZ, 85287-1604; DAVID A. HOSTUTLER, Department of Chemistry, University of Kentucky, Lexington, Kentucky 40506-0055.

Iron monocarbide is the most thoroughly studied transition metal carbide. Here, we report on the study of numerous branch features in the (0,0) [12.0]Ω = 2 ← X^{3}Δ_{3} and (0,0) [13.1]^{3}Φ_{4} ← X^{3}Δ_{3} band systems using high-resolution optical Stark spectroscopy. These intense band systems have been characterized very recently.^{1,2} The electric field induced splitting was analyzed to produce values for the permanent electric dipole moments, μ, of 4.02(6) D, 4.44(6) D and 2.36(3) D for the [12.0]Ω = 2, [13.1]^{3}Φ_{4}, and X^{3}Δ_{3} states, respectively. A comparison with other iron containing diatomic molecules and theoretical predicted values is made. The recent finite field ab initio prediction^{3} of μ is in reasonable agreement with the value of μ(X^{3}Δ). Earlier calculations,^{4} which employed an expectation value approach, are in poor agreement with the current experiment.

1. J.W.-H. Leung, W.S. Tam, Q. Ran, and A.S-C. Cheung, Chem. Phys. Lett. 343, 64 (2001).

2. M. Fujitake, A. Toba, M. Mori, F. Miyazawa, N. Ohashi, K. Aiuchi, and K. Shibuya, J. Mol. Spectrosc. 208, 253 (2001).

3. S. S. Itono, T. Taketsugu, T. Hirano and U. Nagashima, J. Chem. Phys. 115, 11213 (2001).

4. I. Shim and K.A. Gingerich, Eur. Phys. J. D 7, 163 (1999).